OpenStudy (anonymous):

Find the perimeter of the given triangle. triangle WZX,if triangle WZXsimiliar to triangle SRT ST=6,WX=5, and the perimeter of triangle SRT=15

hero (hero):

Perimeter of WZX = 12.5

OpenStudy (anonymous):

ok can you show me how?

hero (hero):

If two triangles are similar, then the ratios of their sides will be equivalent. Which means the ratio of their perimeters will be equal to that ratio. The ratio of ST/WX = 1.2 , Therefore, the ratio of the perimeters will also be 1.2. Knowing all of this, we can set up a proportion to find out the perimeter of the WZX: ST/WX = SRT/WZX 6/5 = 15/x 6x = 75 x = 75/6 x = 12.5

OpenStudy (anonymous):

thanks